A+B+C=360degree cotA/4+cotB/4+cotC/4= kcotA/4cotB/4cotC/4 then k=

Grade:9

1 Answers

sai sandeep
6 years ago
when A+B+C=90 or 360 
A+B+C=90
A+B=90-C
cot(A+B)=Cot(90-C)
cotAcotB-1=cotA+cotB/cotC
cotAcotBcotC-cotC=cotA+cotB
cotA+cotB+cotC=cotAcotBcotC
cotA+cotB+cotC=cotAcotBcotC  
cotA/4+cotB/4+cotC/4=kcotA/4cotB/4cotC/4  
k=1